OpenStudy (anonymous):

Find the area of the regular polygon. Give the answer to the nearest tenth. pentagon with a side of 6 m

OpenStudy (p0sitr0n):

A = 5 triangles A = (b*h)/2 b = 6 h =? <54 degrees sina/a=sinb/b sin72/6 = sin54 /b b = 5.1 h^2 + b^2 = c^2 36 - 5.1^2 = h^2 h = 3.154 A = 6 *3.154 / 2 * 5 = 47.315